OpenStudy (anonymous):

I'm really bad at converting can someone please help me. Convert function into intercept form. y=-x^2+5x+36

OpenStudy (haseeb96):

Would u like to convert the above question in slope intercept form

OpenStudy (anonymous):

No it has to be in \[ y=a(x-r _{1})(x-2_{2})\]

OpenStudy (haseeb96):

Yes u r right.

OpenStudy (haseeb96):

Just use factorizing method in the right side equation and get the this form After factorization.

OpenStudy (haseeb96):

First multiply both sides by - okay? Then it will be -y=x^2-5x-36

OpenStudy (anonymous):

okay but what do you mean by factoring?

OpenStudy (haseeb96):

This method is called factorization method which is given below =x^2-9x+4x-36 -y -y=x(x-9)+4(x-9) -y=(x-9)(x+4)

OpenStudy (haseeb96):

Now again multiply both sides by - in order to make our equation like your given value.

OpenStudy (haseeb96):

My mom is shouting on me.may I go ?

OpenStudy (anonymous):

Okay thanks you for your help

OpenStudy (haseeb96):

So the equation will be y=-(x-9)(x+4) this is your answer.

OpenStudy (anonymous):

Thank you so much(:

OpenStudy (haseeb96):

Just help the people GOD will help u and u feel pleasure after help somebody....welcome my sister.